# Milky Way Globular Clusters: Nurseries for Dynamically-Formed Binary Black Holes

### v1

@@ 3 Results
Here we present the main results of our study. In Section 3.1 we examine the formation and destruction of GCs and their properties in terms of age and metallicity. In Section 3.2 we explore the mass distribution of dynamically-formed BBHs originating within GCs, with particular attention to their birth environments (Sec. 3.3). In Section 3.4 we report the BBH birth and merger rate densities as a function of redshift. Finally, in Section 4 we compare our predictions with previous studies.
3.1 Properties of GCs and comparison with observations
@@
In this section we present the physical properties of GCs in our simulations, including their distributions in mass, metallicity, and age. We compare these characteristics to current observational data for MW GCs (Kruijssen+2015).
@@
Figure 1: (a) Global birth rate density of GCs in the GAMESH simulation as a function of redshift (green solid line). Black lines represent GCs that will not reach z=0 due to either galaxy mergers (dashed line) or gas interactions (dotted line), whereas blue lines show GCs that survive either in the MW halo (solid) or in the disk (dashed) until today; (b) Age-metallicity distribution of surviving MW GCs formed in-situ in the MW galaxy main progenitors (blue dots) and accreted through galaxy mergers (blue stars). Red dots with error bars correspond to individual GCs observed in our Galaxy (Kruijssen+2015). Conversely, green dots represent all GC formation environments within our LG-like volume.
@@
We begin by examining the GC birth rate density within our LG-like volume, as shown by the green solid line in Fig 1(a). Our simulation distinguishes among three GC populations based on their fate: those destroyed by galactic merger events, those destroyed by interactions with giant molecular clouds, and those that survive to $z=0$. As expected, excluding these destruction mechanisms would significantly overestimate the number of surviving GCs. The surviving population is further categorized by location, as some GCs remain in the MW halo and others reside in its disk.
@@
Interestingly, our simulation reveals a peak in the total GC formation at slightly higher redshift (z $\sim 4.5$) compared to previous studies (z $\sim 3$; Bruel+2024; DeLucia+2024) that focus on cosmologic boxes. [^2] Bruel+2024 use the FIRE simulation, which evolves a cube volume of (22.1cMpc)3, while DeLucia+2024 consider a fraction (about 10%) of the entire volume of the Millennium Simulation (Springel+2005). This earlier peak likely stems from the overdense nature of our simulated region, which promotes an earlier onset of both star and GC formation than in average-density environments.
@@
The surviving GC population in the MW halo exhibits a multi-peaked distribution. The peaks are a direct result of major mergers with Lyman-$\alpha$ cooling halos, which efficiently bring a large number of GCs into the MW halo. Conversely, the smooth trend that connects these peaks is primarily driven by continuous, minor mergers with mini-halos. Despite the apparent misalignment with the single formation peak inferred from observed GCs in the MW halo (at $z\sim 3$), our results are in broad agreement with observations when considering the age uncertainties. The typical error bars on observed GC ages, ranging from 0.25 to 0.50 Gyr, are large enough to make our simulated peaks consistent with the single observed peak. The dominant mechanisms for GC destruction also show a clear redshift dependence. At high redshift (z ¿ 10), the cruel cradle effect is the primary driver of disruption. The compact nature and smaller physical sizes of early galaxies result in frequent and highly disruptive interactions with giant molecular clouds. In contrast, at 1 ¡ z ¡ 10, the influence of galaxy mergers surpasses that of gas interactions, becoming the leading cause of GC disruption. As redshift decreases towards z $\sim 1$, a notable fraction of GCs begin to survive within the disk of the MW main progenitor. This trend reflects the declining galaxy’s merger rate and the efficiency of the cruel cradle effect, which is driven by a significant decrease in the interstellar medium (ISM) density within the galaxy’s half-mass radius at later times. Figure 1(b) presents the age-metallicity distribution of GCs surviving until redshift z = 0 in the halo of our MW-like galaxy. They exhibit a significant agreement with the estimated lifetimes of observed GCs. Nevertheless, three of the youngest observed objects deviate from our predictions. These GCs reside in a redshift range where major mergers, typically responsible for introducing GCs into the MW halo, no longer occur (z ¡ 1). It is crucial to acknowledge, however, that the substantial uncertainties in current GC age estimates preclude a conclusive determination of their true formation times. Indeed, taking into account the error bars, two of these GCs overlap with the region of our surviving GCs. Our model struggles to reproduce the oldest metal rich GCs. This is because the oldest GCs are generally expected to have low metallicities, reflecting the pristine conditions of the early Universe. The metallicities of these objects, instead, are comparable to formation environments typically found at lower redshifts (cluster age $\sim 8\div 9$ Gyrs), which is challenging to explain within our current framework. Our simulated GCs tend to have systematically higher values of the metallicity ($\sim 10^{-1}$ Z), approximately one order of magnitude higher than the bulk of observed GCs ($\sim 3\,\times 10^{-2}$ Z). This discrepancy is evident when comparing the simulated data (blue markers) with the observational data in Fig. 1(b). A direct quantitative comparison of metallicity is not straightforward. Observational studies typically report metallicity as [Fe/H], whereas the GAMESH simulation defines metallicity as the abundance of elements heavier than helium (Graziani+2015). This fundamental difference in the definition of metallicity likely contributes to the offset between our simulation results and observations. Beyond this difference, the offset could also be influenced by how metals are diluted within the ISM in our simulation. Specifically, our model assumes that supernova-driven outflows eject gas and metals in proportions identical to the ISM, thus not altering ISM metallicity. However, these outflows might actually be more metal rich then the surrounding ISM. Furthermore, our simulation assumes homogeneous accretion of gas at the average intergalactic medium metallicity. On the other hand, enrichment processes might be far more inhomogeneous, leading to a preferential accretion of unenriched gas, which would dilute the ISM metallicity. Understanding the origin of the MW GCs is another important aspect. We simultaneously present on the same Fig. 1(b) all GC birth environments (green dots), systems which satisfy the gas surface density threshold and have a non-zero cluster formation efficiency. This reveals a wide range of environments, including those that are metal poor (Z $\sim 10^{-2}$ Z). We observe that the majority of survived GCs (blue dots) follow a well-defined metallicity pattern consistent with their birth in the MW main progenitor. Conversely, those deviating from this path originate in satellite galaxies that later merged with the MW. Their distinct metallicity distributions provide a clear way to distinguish accreted GCs from those native to our Galaxy. Additional properties of the survived GCs, including their mass-metallicity distribution, are presented in Appendix A. This analysis reveals that our GC sample is representative of the entire spectrum of observed masses.
3.2 Population of detectable dynamically-formed BBHs
@@
Figure 2: Primary (M1) and secondary (M2) BH mass distribution of observable dynamically-formed BBHs within our LG-like volume. Red dots show the estimated median values of the most massive GW events ($_{1}>50 M$) detected by the LVK collaboration up to O4a, with error bars representing the 90% credible intervals. Dashed lines in histograms display the mass distributions of all BBH mergers that occur in our simulation.
@@
In this section we provide an in-depth analysis of dynamically-formed BBH mergers during our LG’s assembly, focusing on those detectable by current terrestrial interferometers. Figure 2 shows the mass distribution of primary (M1) and secondary (M2) BHs of dynamically-formed BBHs in GCs. The color code in the density map represents the volume density of BBH mergers in our LG. Brighter (darker) regions indicate higher (lower) event occurrence densities. To assess their detectability, we adopt GWFish (Dupletsa+2023), a dedicated Python package based on Fisher matrices for GW population studies. We set a signal-to-noise ratio threshold of 8 to identify events that would be observable by the current network of LVK detectors (GWTC4v2+2025). The primary BH mass distribution offers several insights into the origin of the LVK detections, because different CPS codes yield distinct predictions for the mass range of dynamically-formed BHs. The FASTCLUSTER code struggles to explain GW events with $_{1}<15 M$ through dynamical formation processes. In contrast, such events are plausible, albeit with lower probability densities, in RAPSTER.
@@
An interesting aspect of the M1 distribution is the presence of two prominent peaks in the distribution predicted by both CPS. The lower-mass peak, predicted by both CPS codes around $_{1}\sim 35 M$, appears also in BPS codes but is significantly less prominent in those models (GiacobboMapelli+2018; Belczynski+2020; Tanikawa+2020; Broekgaarden+2022; Iorio+2023). The key difference, indicative of the hierarchical formation scenario, is the presence of the second peak ($_{1}\sim 70-90 M$) which is a distinguishing feature of the CPS predictions. The presence of MBBHs (e.g., $_{1}>50 M$) within the predictions of both codes strongly suggests a dynamical origin for these massive systems. A tail of high-mass mergers extends even into the regime of IMBHs ($_{BH}\sim 10^{2}\div 10^{4} M$) in the RAPSTER simulations, but not in FASTCLUSTER. Despite the formation of thousands of IMBHs in our RAPSTER simulations, just tens of them are potentially detectable by current GW observatories with very low signal-to-noise ratio. This is due to the high masses and merger redshift of these systems, which will be shown in detail in Section 3.4.
Figure 3: (a) Dark matter mass (MDM) and stellar mass (M) distribution of halos that host GCs capable of forming MBBHs; (b) GC mass (MCluster) versus absolute metallicity of GCs hosting MBBHs. The color of each point represents the central stellar number density of the cluster.
@@
The location of IMBHs is an important aspect that our approach is able to provide. Due to their high masses, GW kicks are not large enough to eject these IMBHs from their host stellar clusters. Most of these clusters do not survive as they are destroyed by interactions with gas or by galaxy mergers. Consequently, the IMBHs that once resided in such clusters become “wandering IMBHs” within galaxies. Galaxy mergers are chaotic dynamical processes and it is challenging to precisely trace the final position of the BHs. however, in the case of GCs destroyed by the cruel cradle effect, it is plausible that their IMBHs remain confined to the disk. Here, due to dynamical friction (Capuzzo2008; ArcaSedda+2014), they will tend to migrate towards the galactic center, potentially contributing to the formation of the super-massive BH in the nucleus. The last column of Table 1 shows the number of IMBHs that remain confined to the MW disk. These BHs can grow through mergers to masses spanning from $10^{3}$ M to $4\times 10^{3}$ M. This mass range represents a lower limit, as our work does not account for any gas accretion processes.
@@
Only a minority of IMBHs form within GCs that survive until z = 0. These IMBHs, remaining bound to their host clusters, could therefore be observable in the halo of our Galaxy. The last column of Table 1 shows, in parenthesis, the number of IMBHs that migrate into the Galaxy’s halo together with their GCs. In the halo, they can grow to a mass range of $\sim 2\times 10^{3}$ M. So, this work provides crucial insights into the expected distribution and potential observational locations of IMBHs, guiding future observational surveys. A more detailed discussion of this topic will be presented in future papers. Table 1 presents the statistics of merging BBHs from the GAMESH simulation. As the table shows, the number of BBH mergers is comparable between the two CPS codes and this consistency persists even for systems observable by terrestrial interferometers. By comparing the first and second column, we see that a significant population of BBHs remains undetectable. This happens for several key reasons, primarily related to the frequency of the GWs produced and the sensitivity of the detectors. A major limiting factor is the specific frequency range of terrestrial detectors like LIGO, Virgo, and KAGRA, which are mostly sensitive from approximately 10 Hz to a few kHz. Many BBH mergers, especially those involving very massive BHs ($_{BH}>100 M$) or those in their early inspiral phase, produce GWs at frequencies well below this sensitivity range.
@@ 3.3 Properties of Galaxies and Globular Clusters hosting Massive Binary Black Hole Formation
Figure 3(a) illustrates the properties of the halos that host GCs capable of forming MBBHs. The log-log plot displays a strong correlation between the dark matter mass (MDM) and the stellar mass (M) of these environments. This relationship, which spans several orders of magnitude, closely follows the well-established stellar mass-halo mass relation observed for galaxies (Moster+2013). The tightness of this correlation suggests that the formation of these specific GCs is not a stochastic process, but it is fundamentally linked to the overall mass and evolutionary state of their host galaxy.
The histograms along the axes provide further insight into the mass distributions. The top histogram shows that the majority of these environments have a dark matter mass between $10^{10}\div 10^{12}$ M, a range that includes also progenitors of the MW. The corresponding stellar mass distribution is shown in the left histogram, and it does not exhibit a significant peak that would suggest a preferred range of values.
Figure 4: (a) Redshift evolution of dynamically-formed BBH merger rate density, as a function of primary BH mass (M1). Red curves show BBH with $_{1}\leq 100 M$, green lines indicate MBBHs with 100100 $_{\odot}< M$$_{1}\leq 1000$ M, and violet histograms illustrate IMBHs with $_{1}>1000 M$. Blue star with error bars represents the value provided by the LVK collaboration (GWTC4v2+2025); (b) Redshift evolution of dynamically-formed BBH birth rate density, as a function of primary BH mass (M1). Red curves show BBH with $_{1}\leq 100 M$, green lines indicate MBBHs with 100100 $_{\odot}< M$$_{1}\leq 1000$ M, and violet histograms illustrate IMBHs with $_{1}>1000 M$.
@@
In Fig. 3(b) we show the absolute metallicity distribution of these GCs as a function of their mass, both on a logarithmic scale. Each point represents an individual cluster, and their color indicates the central stellar number density. Unlike the strong correlation in the previous plot, there is no clear relationship between the cluster mass and metallicity. Most GCs have metallicities in the range of $10^{-3}\div 10^{-2}$ and a mass typically in the upper tail of our initial distribution, i.e. between $10^{6}\div 10^{7}$ M. Higher density clusters (yellow points) tend to populate the less massive regions of the plot, while lower density clusters (green points) are found closer to the upper mass limit. This suggests that the central stellar number density and the total mass of a GC play complementary roles in the formation of massive BHs. A lower central density can be compensated for by a higher mass, whereas less massive GCs require a higher core density to form MBBHs. However, the plot indicates a necessary lower limit for central density ($10^{6}$ pc-3) to efficiently form massive BHs through hierarchical mergers.
Figure 3(b) shows that GCs forming MBBHs are generally the most massive and densest clusters that we were able to extract starting from our initial power law distribution (Sec. 2.2). This correlation is significant because it connects the two key requirements for MBBH formation. Indeed, in a high-density environment, stars and BHs are much closer together. This dramatically increases the rate of gravitational encounters and binary interactions. Moreover, the most massive clusters are the ones with the largest potential reservoirs of stellar-mass BHs. So, the high density in their cores provides the dynamic environment needed to efficiently merge these BHs into a more massive object.
The presence of massive, high-density clusters in our sample directly supports the hypothesis that the dynamical conditions necessary for forming MBBHs via hierarchical mergers are present in our simulation.

### v2

@@ 3 Results
Here we present the main results of our study. In Section 3.1 we examine the formation and destruction of GCs and their properties in terms of age and metallicity. In Section 3.2 we explore the mass distribution of dynamically formed BBHs originating within GCs, with particular attention to their birth environments (Sec. 3.3). In Section 3.4 we report the BBH birth and merger rate densities as a function of redshift. Finally, in Section 4 we compare our predictions with previous studies.
3.1 Properties of globular clusters and comparison with observations
@@
In this section we present the physical properties of GCs in our simulations, including their distributions in mass, metallicity, and age. We compare these characteristics to current observational data for MW GCs (Kruijssen 2015).
@@
Figure 1: (a) Global birth rate density of GCs in the GAMESH simulation as a function of redshift (solid green line). Black lines represent GCs that will not reach z=0 due to either galaxy mergers (dashed line) or gas interactions (dotted line), whereas blue lines show GCs that survive either in the MW halo (solid) or in the disk (dashed) until today. (b) Age-metallicity distribution of surviving MW GCs formed in situ in the MW galaxy main progenitors (blue dots) and accreted through galaxy mergers (blue stars). Red dots with error bars correspond to individual GCs observed in our Galaxy (Kruijssen 2015). Conversely, green dots represent all GC formation environments within our LG-like volume.
@@
We begin by examining the GC birth rate density within our LG-like volume, as shown by the solid green line in Fig 1(a). Our simulation distinguishes among three GC populations based on their fate: those destroyed by galactic merger events, those destroyed by interactions with giant molecular clouds, and those that survive to $z=0$. As expected, excluding these destruction mechanisms would significantly overestimate the number of surviving GCs. The surviving population is further categorized by location, as some GCs remain in the MW halo and others reside in its disk.
@@
Interestingly, our simulation reveals a peak in the total GC formation at slightly higher redshift (z $\sim 4.5$) compared to previous studies (z $\sim 3$; Bruel et al. 2024; De Lucia et al. 2024) that focus on cosmologic boxes. [^2] Bruel et al. (2024) use the FIRE simulation, which evolves a cube volume of (22.1 cMpc)3, while De Lucia et al. (2024) consider a fraction (about 10%) of the entire volume of the Millennium Simulation (Springel et al. 2005). This earlier peak likely stems from the overdense nature of our simulated region, which promotes an earlier onset of both star and GC formation than in average-density environments.
@@
The surviving GC population in the MW halo exhibits a multi-peaked distribution. The peaks are a direct result of major mergers with Lyman-$\alpha$ cooling halos, which efficiently bring a large number of GCs into the MW halo. Conversely, the smooth trend that connects these peaks is primarily driven by continuous, minor mergers with mini-halos. Despite the apparent misalignment with the single formation peak inferred from observed GCs in the MW halo (at $z\sim 3$), our results are in broad agreement with observations when considering the age uncertainties. The typical error bars on observed GC ages, ranging from 0.25 to 0.50 Gyr, are large enough to make our simulated peaks consistent with the single observed peak. The dominant mechanisms for GC destruction also show a clear redshift dependence. At high redshift (z ¿ 10), the cruel cradle effect is the primary driver of disruption. The compact nature and smaller physical sizes of early galaxies result in frequent and highly disruptive interactions with giant molecular clouds. In contrast, at 1 ¡ z ¡ 10, the influence of galaxy mergers surpasses that of gas interactions, becoming the leading cause of GC disruption. As redshift decreases toward z $\sim 1$, a notable fraction of GCs begin to survive within the disk of the MW main progenitor. This trend reflects the decline in the merger rate and in efficiency of the cruel cradle effect, which is driven by a significant decrease in the interstellar medium (ISM) density within the galaxy’s half-mass radius at later times. The sharp decrease in the fraction of disrupted GCs at z $\approx$ 1 is primarily driven by the transition of the MW progenitors – the main contributors to GC formation in our LG-like volume – into a secular evolution phase, where the absence of major mergers and the post-reionization suppression of star formation in satellites significantly reduces the efficiency of GC destruction. Figure 1(b) presents the age-metallicity distribution of GCs surviving until redshift z = 0 in the halo of our MW-like galaxy. They exhibit a significant agreement with the estimated lifetimes of observed GCs. Nevertheless, three of the youngest observed objects deviate from our predictions. On the basis of the age determination, these GCs are predicted to form at z ¡ 1 when major mergers - which are responsible for introducing GCs into the MW halo - no longer occur in our simulation. It is crucial to acknowledge, however, that the substantial uncertainties in current GC age estimates preclude a conclusive determination of their true formation times. Indeed, taking into account the error bars, two of these GCs overlap with the region of our surviving GCs. Our model struggles to reproduce the oldest metal rich GCs. This is because the oldest GCs are generally expected to have low metallicities, reflecting the pristine conditions of the early Universe. The metallicities of these objects, instead, are comparable to formation environments typically found at lower redshifts (cluster age $\sim 8-9$ Gyrs), which is challenging to explain within our current framework. Our simulated GCs tend to have systematically higher values of the metallicity ($\sim 10^{-1}$ Z), approximately one order of magnitude higher than the bulk of observed GCs ($\sim 3\,\times 10^{-2}$ Z). This discrepancy is evident when comparing the simulated data (represented in blue) with the observational data in Fig. 1(b). A direct quantitative comparison of metallicity is not straightforward. Observational studies typically report metallicity as [Fe/H], whereas the GAMESH simulation defines metallicity as the abundance of elements heavier than helium (Graziani et al. 2015). This fundamental difference in the definition of metallicity likely contributes to the offset between our simulation results and observations. Beyond this difference, the offset could also be influenced by how metals are diluted within the ISM in our simulation. Specifically, our model assumes that supernova-driven outflows eject gas and metals in proportions identical to the ISM, thus not altering ISM metallicity. However, these outflows might actually be more metal rich than the surrounding ISM. Furthermore, our simulation assumes homogeneous accretion of gas at the average intergalactic medium metallicity. On the other hand, enrichment processes might be far more inhomogeneous, leading to a preferential accretion of unenriched gas, which would dilute the ISM metallicity. Understanding the origin of the MW GCs is another important aspect. We simultaneously present on the same Fig. 1(b) all GC birth environments (green dots), systems that satisfy the gas surface density threshold and have a nonzero cluster formation efficiency. This reveals a wide range of environments, including those that are metal poor (Z $\sim 10^{-2}$ Z). We observe that the majority of survived GCs (blue dots) follow a well-defined metallicity pattern consistent with their birth in the MW main progenitor. Conversely, those deviating from this path originate in satellite galaxies that later merged with the MW. Their distinct metallicity distributions provide a clear way to distinguish accreted GCs from those native to our Galaxy. Furthermore, there is an interesting population of newly formed GCs with metallicity below 0.01 Z. This population is born within dwarf galaxies located at the periphery of our LG-like volume. These smaller systems, being more isolated and having lower star formation efficiencies, remain more chemically pristine compared to the main hosts. This environmental diversity reflects the bimodal metallicity distribution observed in the plot. The gap between these two populations is also sensitive to the adopted minimum gas surface density threshold for GC formation of $10^{3}$ M pc-2: removing this threshold would indeed reveal lower-metallicity environments, as shown in Appendix C.
Additional properties of the survived GCs, including their mass-metallicity distribution, are presented in Appendix A. This analysis reveals that our GC sample is representative of the entire spectrum of observed masses.
3.2 Population of detectable dynamically formed binary black holes
@@
Figure 2: Primary (M1) and secondary (M2) BH mass distribution of observable dynamically formed BBHs within our LG-like volume. Red dots show the estimated median values of the most massive GW events ($_{1}>50 M$) detected by the LVK collaboration up to O4a, with error bars representing the 90% credible intervals. Dashed lines in histograms display the mass distributions of all BBH mergers that occur in our simulation.
@@
In this section we provide an in-depth analysis of dynamically formed BBH mergers during our LG’s assembly, focusing on those detectable by current terrestrial interferometers. Figure 2 shows the mass distribution of primary (M1) and secondary (M2) BHs of dynamically formed BBHs in GCs. The color code in the density map represents the volume density of BBH mergers in our LG. Brighter (darker) regions indicate higher (lower) event occurrence densities. To assess their detectability, we adopt GWFish (Dupletsa et al. 2023), a dedicated Python package based on Fisher matrices for GW population studies. We set a signal-to-noise ratio threshold of 8 to identify events that would be observable by the current network of LVK detectors (The LIGO Scientific Collaboration et al. 2025a). The primary BH mass distribution offers several insights into the origin of the LVK detections, because different CPS codes yield distinct predictions for the mass range of dynamically formed BHs. The FASTCLUSTER code struggles to explain GW events with $_{1}<15 M$ through dynamical formation processes. In contrast, such events are plausible, albeit with lower probability densities, in RAPSTER.
@@
An interesting aspect of the M1 distribution is the presence of two prominent peaks in the distribution predicted by both CPS. The lower-mass peak, predicted by both CPS codes around $_{1}\sim 35 M$, appears also in BPS codes but is significantly less prominent in those models (Giacobbo et al. 2018; Belczynski et al. 2020; Tanikawa et al. 2020; Broekgaarden et al. 2022; Iorio et al. 2023). The key difference, indicative of the hierarchical formation scenario, is the presence of the second peak ($_{1}\sim 70-90 M$) that is a distinguishing feature of the CPS predictions. The presence of MBBHs (e.g., $_{1}>50 M$) within the predictions of both codes strongly suggests a dynamical origin for these massive systems. A tail of high-mass mergers extends even into the regime of IMBHs ($_{\rm BH}\sim 10^{2}-10^{4} M$) in the RAPSTER simulations, but not in FASTCLUSTER. Despite the formation of thousands of IMBHs in our RAPSTER simulations, just tens of them are potentially detectable by current GW observatories with very low signal-to-noise ratio. This is due to the high masses and merger redshift of these systems, which will be shown in detail in Section 3.4.
Figure 3: (a) Dark matter mass (MDM) and stellar mass (M) distribution of halos that host GCs capable of forming MBBHs. (b) Two-dimensional density distribution of GCs in the mass-metallicity plane. The contours represent the density estimation of the cluster population, highlighting the regions where initial central stellar number density and total mass facilitate MBBH formation, with color gradients distinguishing between high initial central number density (red-yellow tones, $\rho_{c}>$ $5\times 10^{7}$ pc-3) and lower initial central density (blue-green tones, $\rho_{c}<$ $5\times 10^{7}$ pc-3).
@@
The location of IMBHs is an important aspect that our approach is able to provide. Due to their high masses, GW kicks are not large enough to eject these IMBHs from their host stellar clusters. Most of these clusters do not survive as they are destroyed by interactions with gas or by galaxy mergers. Consequently, the IMBHs that once resided in such clusters become ”wandering IMBHs” within galaxies. Galaxy mergers are chaotic dynamical processes and it is challenging to precisely trace the final position of the BHs. however, in the case of GCs destroyed by the cruel cradle effect, it is plausible that their IMBHs remain confined to the disk. Here, due to dynamical friction (Capuzzo-Dolcetta and Miocchi 2008; Arca-Sedda and Capuzzo-Dolcetta 2014), they will tend to migrate toward the galactic center, potentially contributing to the formation of the super-massive BH in the nucleus. The last column of Table 1 shows the number of IMBHs that remain confined to the MW disk. These BHs can grow through mergers to masses spanning from $10^{3}$ M to $4\times 10^{3}$ M. This mass range represents a lower limit, as our work does not account for any gas accretion processes.
@@
Only a minority of IMBHs form within GCs that survive until z = 0. These IMBHs, remaining bound to their host clusters, could therefore be observable in the halo of our Galaxy. The last column of Table 1 shows, in parenthesis, the number of IMBHs that migrate into the Galaxy’s halo together with their GCs. In the halo, they can grow to a mass range of $\sim 2\times 10^{3}$ M. So, this work provides crucial insights into the expected distribution and potential observational locations of IMBHs, guiding future observational surveys. A more detailed discussion of this topic will be presented in future papers. Table 1 presents the statistics of merging BBHs from the GAMESH simulation. As the table shows, the number of BBH mergers is comparable between the two CPS codes and this consistency persists even for systems observable by terrestrial interferometers. By comparing the first and second column, we see that a significant population of BBHs remains undetectable. This happens for several key reasons, primarily related to the frequency of the GWs produced and the sensitivity of the detectors. A major limiting factor is the specific frequency range of terrestrial detectors, such as LIGO, Virgo, and KAGRA, which are mostly sensitive from approximately 10 Hz to a few kHz. Many BBH mergers, especially those involving very massive BHs ($_{\rm BH}>100 M$) or those in their early inspiral phase, produce GWs at frequencies well below this sensitivity range.
@@ 3.3 Properties of galaxies and globular clusters hosting massive binary black hole formation
Figure 3(a) illustrates the properties of the halos that host GCs capable of forming MBBHs. The log-log plot displays a strong correlation between the dark matter mass (MDM) and the stellar mass (M) of these environments. This relationship, which spans several orders of magnitude, closely follows the well-established stellar mass-halo mass relation observed for galaxies (Moster et al. 2013). The tightness of this correlation suggests that the formation of these specific GCs is not a stochastic process, but it is fundamentally linked to the overall mass and evolutionary state of their host galaxy.
The histograms along the axes provide further insight into the mass distributions. The top histogram shows that the majority of these environments have a dark matter mass between $10^{10}-10^{12}$ M, a range that includes also progenitors of the MW. The corresponding stellar mass distribution is shown in the left histogram, and it does not exhibit a significant peak that would suggest a preferred range of values.
Figure 4: (a) Redshift evolution of dynamically formed BBH merger rate density, as a function of primary BH mass (M1). Red curves show BBH with $_{1}\leq 100 M$, green lines indicate MBBHs with $100$ $_{\odot}< M$$_{1}\leq 1000$ M, and violet histograms illustrate IMBHs with $_{1}>1000 M$. Blue star with error bars represents the value provided by the LVK collaboration (The LIGO Scientific Collaboration et al. 2025a). (b) Redshift evolution of dynamically formed BBH birth rate density, as a function of primary BH mass (M1). Red curves show BBH with $_{1}\leq 100 M$, green lines indicate MBBHs with $100$ $_{\odot}< M$$_{1}\leq 1000$ M, and violet histograms illustrate IMBHs with $_{1}>1000 M$.
@@
In Fig. 3(b) we show the absolute metallicity distribution of these GCs as a function of their mass, both on a logarithmic scale. The 2D contours represent the number density distribution of clusters, while the marginal histograms on the top and left axes show the 1D distributions for mass and metallicity, respectively. The color gradient indicates the initial central stellar number density ($\rho_{c}$): red-yellow tones represent higher density clusters ($\rho_{c}>$ $5\times 10^{7}$ pc-3), while blue-green tones correspond to lower density ones ($\rho_{c}\leq$ $5\times 10^{7}$ pc-3).
Most GCs have metallicities in the range of $10^{-3}-10^{-2}$ and a mass typically in the upper tail of our initial distribution, i.e., between $10^{6}-10^{7}$ M. Higher density clusters (red-yellow tones) tend to populate the less massive regions of the plot, while lower density clusters (blue-green tones) are found closer to the upper mass limit. This suggests that the initial central stellar number density and the total mass of a GC play complementary roles in the formation of massive BHs. A lower central density can be compensated for by a higher mass, whereas less massive GCs require a higher core density to form MBBHs. However, the plot indicates a necessary lower limit for initial central density (¿ $10^{6}$ pc-3) to efficiently form massive BHs through hierarchical mergers.
Figure 3(b) shows that GCs forming MBBHs are generally the most massive and densest clusters that we were able to extract starting from our initial power law distribution (Sec. 2.2). This correlation is significant because it connects the two key requirements for MBBH formation. Indeed, in a high-density environment, stars and BHs are much closer together. This dramatically increases the rate of gravitational encounters and binary interactions. Moreover, the most massive clusters are the ones with the largest potential reservoirs of stellar-mass BHs. So, the high density in their cores provides the dynamic environment needed to efficiently merge these BHs into a more massive object. Consequently, the presence of massive, high-density clusters in our sample directly supports the hypothesis that the dynamical conditions necessary for forming MBBHs via hierarchical mergers are present in our simulation.